A parking lot has cars and motorcycles . altogether there are 18 vechicles and 66 wheels . which system of linear equations represents the problem. let m be the number of motorcycles and c be the number of cars
Answer by Edwin McCravy(20054)

A parking lot has cars and motorcycles . altogether there are 18 vechicles and 66 wheels . which system pf linear equations represents the problem. let m be the number of motorcycles and c be the number of cars
  m + c = 18
2m + 4c = 66


But you can do the problem in your head:

If they were all motocycles there would be 36 wheels.
But there are 66 wheels, so the remaining 30 wheels
belong to cars. 30 wheels is 15 pairs of wheels.
So there are 15 cars and 18-15 = 3 motorcycles.

Edwin